OpenStudy (anonymous):

The length of a rectangle is 5 meters more than its width. If the area is 66 square meters, what are the length and width? Show all Work!

OpenStudy (anonymous):

Let the width be x then length will be x + 5 Area = x(x + 5) given area = 66 square meters x(x + 5) = 66 x^2 + 5x = 66 x^2 + 5x - 66 = 0 (x + 11)(x - 6) = 0 x = -11, 6 ignore negative value x = 6 width = 6 m length = 6 + 5 = 11 m

OpenStudy (anonymous):

Let the width be x then length will be x + 5 Area = x(x + 5) given area = 66 square meters x(x + 5) = 66 x^2 + 5x = 66 x^2 + 5x - 66 = 0 (x + 11)(x - 6) = 0 x = -11, 6 ignore negative value x = 6 width = 6 m length = 6 + 5 = 11 m

OpenStudy (radar):

Let w=width Then length = w+5 meters Given Area is 66 sq meters. Area is equal to length times the width and is 66 sq meters (w+5) w = 66 w^2+5w =66 w^2 + 5w -66=0 (standard form) factor (w+11)(w-6)=0 w+11=0, w=-11, w-6=0, w=6 Use the positive value w=6 and length equal w+5 or length =11 Just as chris also worked out above, that was a good answer.

OpenStudy (anonymous):

thanks:-D

OpenStudy (anonymous):

Thank you all